OpenStudy (anonymous):

A number cube with the numbers 1, 2, 3, 4, 5, and 6 is rolled 1000 times. About how many times would it be expected that a number greater than 4 is rolled? A. 667 B. 167 C. 500 D. 333

OpenStudy (anonymous):

@crazysingh

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@CO_oLBoY

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

3/6 = 1/2 chance of rolling 4 or above so if you roll a cube 1000 times the chances of you making 4 or above is half so the answer is 500

OpenStudy (anonymous):

no the ans is 333

OpenStudy (anonymous):

how

OpenStudy (anonymous):

coz greater than 4 are 2 digits

OpenStudy (anonymous):

ranga uncle what do u say

OpenStudy (anonymous):

4 or greater meaning 4,5,6

OpenStudy (anonymous):

so (2/6)x1000 =333

OpenStudy (ranga):

Greater than 4 here is getting a 5 or a 6. P(5) = 1/6 P(6) = 1/6 P(5) or P(6) = P(5) + P(6) = 1/6 + 1/6 = 2/6 = 1/3 If rolled 1000 times, the number of times of getting either a 5 or a 6 is 1000 * 1/3 = 333.

OpenStudy (anonymous):

out of six times we can expect number greater than four = 2 times out of one time we can expect number greater than four = 1/3 times out of 1000 times we can expect number greater than four = 1000/3= 333.33 times

OpenStudy (anonymous):

no it says 'greater than 4' not '4 or geater than it'

OpenStudy (anonymous):

the question isn't asking greater than 4 the question is asking 4 OR above

OpenStudy (anonymous):

@K1NG

OpenStudy (anonymous):

@K1NG read the question again